# Living with Data: Exploring Physicalization Approaches to Sedentary Behavior Intervention for Older Adults in Everyday Life

### v2

@@ Abstract.
Sedentary behavior is a critical health risk for older adults. Although digital interventions are widely available, they primarily rely on screen-based notifications that can feel clinical or cognitively demanding, and are thus often ignored over time. This paper presents a three phase Research through Design methodology to explore data physicalization approaches that ambiently represented sedentary data patterns using decor artifacts in older adults’ homes. These artifacts transformed abstract data into aesthetic, evolving forms, that became part of the domestic landscape. Our research revealed how these physicalizations fostered self-reflection, family conversations, and encouraged active lifestyles. We demonstrate how qualities like aesthetic ambiguity and slow revelation can empower older adults, fostering a reflective relationship with their well-being. Ultimately, we argue that creating data physicalizations for older adults necessitates a shift from merely informing users to enabling them to live with, and through, their data.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. The CushLumi system embedded in an older adult’s living room: flexible sensors inside a familiar seat cushion capture sedentary behavior and send sensor data to a cloud-based processor, which in turn drives an ambient lamp that gradually changes its light to physicalize sitting patterns. Rather than issuing screen-based alerts, the system weaves personal sedentary data into the domestic atmosphere, gently inviting reflection and movement within everyday life.:
@@ 1. Introduction
A suite of health and wellness technologies has been developed to help people monitor their daily activity, sleep, and sedentary behavior (gao2019emerging; sugiyama2008joint; owen2010sedentary). These technologies include wearable trackers and smartwatches that visualize one’s steps, heart rate, and sitting time via dashboards and notifications (Reinschluessel2025; tremblay2010physiological). They also include home-based systems that provide ambient cues and persuasive prompts, such as light- or and sound-based reminders to encourage movement breaks (10.1145/3640794.3665561; 7927391; Yu2025). Such systems have been designed for contexts that value numerical precision and individual self optimization, emphasizing screen-based feedback, charts, and goals as primary interaction paradigms (info:doi/10.2196/54375; info:doi/10.2196/25171).
However, with the aging population and the widespread diffusion of consumer health devices, these systems are located in households where needs and norms often differ from their original design contexts. Older adults, for example, often prioritize simplicity, comfort, and emotional ease over granular analytics (app15105458). The home is also a social and aesthetic setting where devices are expected to blend in with furniture, routines, and rhythms rather than demand active attention (Ghorayeb2023; Yu2025). This prompts reflection on the suitability of screen- and number-centric tools for people who may have increased cognitive load, fine motor challenges, or anxiety from constant alerts (Sagga2025; Miles2024). For sedentary behavior in particular, the gap between raw sensor readings and immediately meaningful cues in everyday life remains underexplored (info:doi/10.2196/54375).
@@
As a result of our research, we make three contributions. First,we extend prior work on data physicalization by understanding the potential and perceptions of data physicalizations in real world settings. These findings were identified via a three-phase Research through Design process that culminated in an in-situ deployment of an ambient data physicalization system (i.e., the CushLumi) in older adults’ homes. Second, we advance the notion of living with data by proposing a perspective shift that allowed for personal sedentary data to be integrated into a physical model that shaped environmental ambiance and felt atmospheres (rather than dashboards or notification alarms). Third, we distill our findings into design implications for future tangible health technologies, highlighting how the decoupling of sensing from display, the use of positive metaphors, and the integration of feedback into familiar domestic objects can preserve the richness of personal histories while offering non-disruptive, experiential feedback that supports older adults’ autonomy and ongoing reflection.
@@ 2. Related Work
A substantial body of research has demonstrated the importance of supporting gentle, situated forms of engagement with everyday information via ambient technologies. Such research has integrated cues into the periphery of daily life, allowing users to remain aware without being interrupted or cognitively overloaded (10.1145/3670653.3670681; 10.1093/iwc/iwae054). In parallel, research on data physicalizations has shown how transforming abstract information into tangible and aesthetic forms can slow interaction, foster curiosity through ambiguity and gradual revelation, and create opportunities for reflective engagement beyond screen-based interfaces (Bae2022; 9200790; miao2025physlens; 10.1145/3706598.3713267). Despite these complementary strengths, relatively little research has explored how ambient qualities can be embedded within data physicalization practices to represent everyday bodily states, such as patterns of sitting or inactivity, or how such hybrid representations might become enduring, meaningful elements within older adults’ domestic environments.
@@
Supporting older adults’ well-being ultimately requires interventions that can foster sustained, long-term changes in everyday activity patterns (Madigan2024; Ominyi2024). However, existing behavior-change technologies have primarily relied on wearables and mobile apps that motivate activity through monitoring, goal setting, and notifications (info:doi/10.2196/54375). While such systems can be effective in the short term, engagement frequently declines over time, limiting their impact on long-term reductions in sedentary behavior (Brickwood2019). More recent studies have begun to explore ambient and embodied prompts that use light, sound, or physical cues to encourage brief standing or walking breaks, demonstrating that situated cues can facilitate micro-behavior changes without disrupting daily routines (10.1145/3640794.3665561; 7927391; Zenscape2024). Many of these interventions tend to be short-term focused and implemented in controlled laboratory settings, making it unclear how well they align with the lived rhythms of everyday environments. This gap underscores the opportunity to create ambient, physicalized representations that sustain interest, invite ongoing interpretation, and gently support adaptive relationships with wellbeing data within the material fabric of everyday domestic life.
@@
To support subtle, long-term behavior change with ambient data physicalizations, it is crucial to understand how older adults prefer everyday behavioral information to be conveyed. Prior research found that older adults valued simplicity, comfort, and low-stress interactions rather than detailed dashboards or frequent alerts (app15105458). Gerontechnology studies further highlighted the importance of minimizing cognitive load and designing technologies that align with the aesthetics and rhythms of domestic life (Laukka2024; Ghorayeb2023). Although existing systems attempt to provide lightweight or non-intrusive visualizations for aging populations (While2024; Blascheck2025), they often remain screen-based or reminder-driven, which is counter to the values of older adults. Considering that furniture and lighting are both influential for older adults’ physical and emotional wellbeing and constitute familiar (Yu2025), pervasive elements of domestic interiors, they offer particularly promising channels for embedding subtle, situated health-related feedback.
@@ 3. Methodology
To understand how sedentary behavior data might be captured, represented, and situated in older adults’ everyday environments, we employed a Research through Design (RtD) methodology (10.1145/1240624.1240704; gaver2012should). Our research had three phases: ethnographic observations (Phase 1), a poster-based discussion with diverse stakeholders (Phase 2), and a prototype deployment (Phase 3). This methodological approach was chosen because our research focuses on a problem that does not have a single optimal solution and requires the generation of new design knowledge through iterative cycles of making and reflecting.
3.1. Participants
For Phase 1, we collaborated with a local nursing home to recruit 15 participants for our ethnographic observations. We then conducted semi-structured interviews with 12 of these participants after the ethnographic observations. For Phase 2, we recruited 9 stakeholders. Three were researchers specializing in geriatric medicine, public health, and biomedical sciences, another three were interaction designers with experience in healthcare and systems for older adults, and the remaining three were older adults who represented our target end users. These stakeholders had diverse perspectives and levels of prior engagement with personal data, clinical, design, and lived-experiences, similar to prior work (9200790). For Phase 3, we recruited 5 older adults with sedentary habits. Each participant hosted the CushLumi systemin their home for a full day and then participated in semi-structured interviews to capture their lived experience using the system. Herein, we refer to older adult participants as P1–P15, expert researchers as R1–R3, and interaction designers as D1–D3 (Table 1).
@@ 3.2. Data Physicalization Design Concepts and the CushLumi Prototype
Figure 2. The schematic diagram of the CushLumi System: A Cloud-Based Smart Home Lighting System Driven by Cushion Sensors. The system consisted three parts: A) a Sensing Cushion and C) an Ambient Light lamp, which are the two physical artefacts users directly engage with, and C) a cloud‑based processing component that acts as an invisible bridge, receiving sit/stand events from the cushion and driving the lamp’s gradual light changes and celebratory animations. Whenever a user sat on the cushion, its deformation was detected and the system began tracking the duration of one’s sitting. This information was then used to change the lamp’s hue and light intensity to motivate the user to stand up and move around.
Insights from Phase 1 led us away from a smartphone‑based intervention toward feedback that could be quietly embedded in everyday surroundings. Participants reported disliking complex apps and “nagging” notifications, and cushions repeatedly appeared as mundane but central artefacts in long sitting episodes. At the same time, prior work highlights cushions and lighting as familiar, pervasive elements of older adults’ domestic interiors and as promising channels for low‑priority ambient feedback (wisneski1998ambient; davis2017effects; mclaughlin2020designing; feine2025integrating). Building on this, we developed CushLumi as a two‑part, hybrid data physicalization system: a sensing cushion that captures sedentary behaviour, and a separate ambient lamp that materializes this data in the home.
autoreffig:schematicdiagram presents an overall schematic of the CushLumi system, while Figure 3, footnote 2 and Figure 5 illustrate its layered architecture, hardware components, and light behaviour. In everyday use, the cushion detects when an older adult sits down, stands up, and remains seated for prolonged periods. These events are processed in a simple cloud‑based pipeline that tracks sit durations and drives a small wooden lamp placed nearby. As uninterrupted sitting accumulates, the lamp’s colour and brightness gradually shift from a soft warm white toward a more intense, reddish glow; when the person stands, the lamp briefly plays a colourful celebratory animation before returning to its baseline state. Laser‑cut patterns on the lamp depict positive everyday activities (e.g., stretching, gardening), and a touch‑sensitive top surface serves as a mute button, allowing users to temporarily silence the feedback and thereby maintain a sense of agency. Further design explorations and rationale via RtD, technical details regarding the sensing hardware, light‑mapping logic, and cloud infrastructure are provided in Appendix A.
@@
Figure 4. The components of the CushLumi system. The CushLumi system major used Particle Photon microcontrollers [^1] https://www.particle.io/, which was programmed in C++, to process real-time data from the flex wiring sensors and control the LED circle. The wireless communication between the hardware components was controlled using Particle’s cloud platform. IFTTT [^2] https://ifttt.com/ was also used for time-based, and remote control, of the system to prototype simple remote and time-based control of the device (e.g., muting or unmuting the light via a phone widget or scheduled applet).
@@ 3.3. Procedures
Herein we describe the procedures for each phase of our RtD methodology.
@@ 3.3.1. Phase 1
To understand the context of sedentary behavior and older adults’ relationship with technology, we first conducted ethnographic observations of 15 older adults in a local nursing home. We focused on their daily routines, sitting habits, and interactions with everyday objects and technologies. Field notes documented locations, durations, and social contexts of sitting, as well as observed uses of digital devices. Following the observations, we conducted semi-structured interviews with 12 of these older adults. The interviews covered topics such as daily activity patterns, attitudes toward sedentary behavior, experiences with digital technologies (including smartphones and other screen-based devices), and preferences for how health-related information should be communicated. Interviews also invited participants to reflect on specific situations observed during the ethnographic phase.
@@ 3.3.3. Phase 3
We obtained informed consent from all participants prior to each phase. For older adults in the nursing home and in-home deployments, the study procedures were also coordinated with facility staff and family members when appropriate. Participants could withdraw at any time without penalty, and all collected data was pseudonymized during analysis and for reporting herein.
@@ 3.4. Data Collection and Analysis
We analyzed the semi-structured interview data using reflexive thematic analysis following Braun and Clarke’s approach, as adapted in prior HCI research (lin2019using; bowman2023using). Two researchers independently open-coded the transcripts, identifying segments related to perceptions of sedentary behavior, reactions to different forms of data representation, emotional responses, and experiences with the prototypes. They then met to discuss and merge their findings into a unified codebook. This codebook was systematically applied to the entire dataset, with the researchers working to refine the codes and group them into higher-level themes. The resulting themes were then related back to the three phases of the RtD process, informing both the iterative design decisions described above and the final set of findings reported herein. Note that we use annotations such as P3_\_Phase1 to indicate that, in this example, a quote was from P3 during Phase 1.

### v3

@@ Abstract.
Sedentary behavior is a critical health risk for older adults. Although digital interventions are widely available, they primarily rely on screen-based notifications that can feel clinical or cognitively demanding, and are thus often ignored over time. This paper presents a three-phase Research through Design methodology to explore data physicalization approaches that ambiently represent sedentary data patterns using decor artifacts in older adults’ homes. These artifacts transformed abstract data into aesthetic, evolving forms that became part of the domestic landscape. Our research revealed how these physicalizations fostered self-reflection, family conversations, and encouraged active lifestyles. We demonstrate how qualities like aesthetic ambiguity and slow revelation can empower older adults, fostering a reflective relationship with their well-being. Ultimately, we argue that creating data physicalizations for older adults necessitates a shift from merely informing users to enabling them to live with and through their data.
@@
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791978isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interaction techniquesccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interaction design process and methodsccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
Figure 1. The CushLumi system embedded in an older adult’s living room: flexible sensors inside a familiar seat cushion capture sedentary behavior and send sensor data to a cloud-based processor, which in turn drives an ambient lamp that gradually changes its light to physicalize sitting patterns. Rather than issuing screen-based alerts, the system weaves personal sedentary data into the domestic atmosphere, gently inviting reflection and movement within everyday life.
the CushLumi embedded in an older adult’s living room: flexible sensors inside a familiar seat cushion capture sedentary behavior and send sensor data to a cloud-based processor, which in turn drives an ambient lamp that gradually changes its light to physicalize sitting patterns. Rather than issuing screen-based alerts, the system weaves personal sedentary data into the domestic atmosphere, gently inviting reflection and movement within everyday life.
Figure 2. The schematic diagram of the CushLumi System: A Cloud-Based Smart Home Lighting System Driven by Cushion Sensors. The system consisted of three parts: A) a Sensing Cushion and C) an Ambient Light lamp, which are the two physical artefacts users directly engage with, and C) a cloud‑based processing component that acts as an invisible bridge, receiving sit/stand events from the cushion and driving the lamp’s gradual light changes and celebratory animations. Whenever a user sat on the cushion, its deformation was detected, and the system began tracking the duration of one’s sitting. This information was then used to change the lamp’s hue and light intensity to motivate the user to stand up and move around.
@@ 1. Introduction
A suite of health and wellness technologies has been developed to help people monitor their daily activity, sleep, and sedentary behavior (Gao et al., 2019; Sugiyama et al., 2008; Owen et al., 2010). These technologies include wearable trackers and smartwatches that visualize one’s steps, heart rate, and sitting time via dashboards and notifications (Reinschluessel et al., 2025; Tremblay et al., 2010). They also include home-based systems that provide ambient cues and persuasive prompts, such as light- or sound-based reminders to encourage movement breaks (Ahire et al., 2024; Davis et al., 2017b; Yu and Juan, 2025). Such systems have been designed for contexts that value numerical precision and individual self-optimization, emphasizing screen-based feedback, charts, and goals as primary interaction paradigms (Zhu et al., 2024; Feng et al., 2021).
However, with the aging population and the widespread diffusion of consumer health devices, these systems are located in households where needs and norms often differ from their original design contexts. Older adults, for example, often prioritize simplicity, comfort, and emotional ease over granular analytics (He et al., 2025). The home is also a social and aesthetic setting where devices are expected to blend in with furniture, routines, and rhythms rather than demand active attention (Ghorayeb et al., 2023; Yu and Juan, 2025). This prompts reflection on the suitability of screen- and number-centric tools for people who may have increased cognitive load, fine motor challenges, or anxiety from constant alerts (Sagga, 2025; Miles et al., 2024). For sedentary behavior in particular, the gap between raw sensor readings and immediately meaningful cues in everyday life remains underexplored (Zhu et al., 2024).
@@
As a result of our research, we make three contributions. First, we extend prior work on data physicalization by understanding the potential and perceptions of data physicalization by evaluating its application specifically for older adults within domestic environments. These findings were identified via a three-phase Research through Design process that culminated in an in-situ deployment of an ambient data physicalization system (i.e., the CushLumi) in older adults’ homes. Second, we operationalize the notion of living with data by proposing a perspective shift that allowed for personal sedentary data to be integrated into a physical model that shaped environmental ambiance and felt atmospheres (rather than dashboards or notification alarms). Third, we distill our findings into design implications for future tangible health technologies, highlighting how the decoupling of sensing from display, the use of positive metaphors, and the integration of feedback into familiar domestic objects can preserve the richness of personal histories while offering non-disruptive, experiential feedback that supports older adults’ autonomy and ongoing reflection.
@@ 2. Related Work
A substantial body of research has demonstrated the importance of supporting gentle, situated forms of engagement with everyday information via ambient technologies. Such research has integrated cues into the periphery of daily life, allowing users to remain aware without being interrupted or cognitively overloaded (Jung-Krenzer et al., 2024; Perera et al., 2024). In parallel, research on data physicalizations has shown how transforming abstract information into tangible and aesthetic forms can slow interaction, foster curiosity through ambiguity and gradual revelation, and create opportunities for reflective engagement beyond screen-based interfaces (Bae et al., 2022; Karyda et al., 2021; Miao et al., 2025; Di Lodovico et al., 2025). Despite these complementary strengths, relatively little research has explored how ambient qualities can be embedded within data physicalization practices to represent everyday bodily states, such as patterns of sitting or inactivity, or how such hybrid representations might become enduring, meaningful elements within older adults’ domestic environments.
@@
Supporting older adults’ well-being ultimately requires interventions that can foster sustained, long-term changes in everyday activity patterns (Madigan et al., 2024; Ominyi et al., 2024). However, existing behavior-change technologies have primarily relied on wearables and mobile apps that motivate activity through monitoring, goal setting, and notifications (Zhu et al., 2024). While such systems can be effective in the short term, engagement frequently declines over time, limiting their impact on long-term reductions in sedentary behavior (Brickwood et al., 2019). More recent studies have begun to explore ambient and embodied prompts that use light, sound, or physical cues to encourage brief standing or walking breaks, demonstrating that situated cues can facilitate micro-behavior changes without disrupting daily routines (Ahire et al., 2024; Davis et al., 2017b; Xie et al., 2024). Many of these interventions tend to be short-term focused and implemented in controlled laboratory settings, making it unclear how well they align with the lived rhythms of everyday environments. This gap underscores the opportunity to create ambient, physicalized representations that sustain interest, invite ongoing interpretation, and gently support adaptive relationships with wellbeing data within the material fabric of everyday domestic life.
@@
To support subtle, long-term behavior change with ambient data physicalizations, it is crucial to understand how older adults prefer everyday behavioral information to be conveyed. Prior research found that older adults valued simplicity, comfort, and low-stress interactions rather than detailed dashboards or frequent alerts (He et al., 2025). Gerontechnology studies further highlighted the importance of minimizing cognitive load and designing technologies that align with the aesthetics and rhythms of domestic life (Laukka et al., 2024; Ghorayeb et al., 2023). Although existing systems attempt to provide lightweight or non-intrusive visualizations for aging populations (While et al., 2024; Blascheck et al., 2025), they often remain screen-based or reminder-driven, which is counter to the values of older adults. Considering that furniture and lighting are both influential for older adults’ physical and emotional well-being and constitute familiar (Yu and Juan, 2025), pervasive elements of domestic interiors, they offer particularly promising channels for embedding subtle, situated health-related feedback.
@@ 3. Methodology
To understand how sedentary behavior data might be captured, represented, and situated in older adults’ everyday environments, we employed a Research through Design (RtD) methodology (Zimmerman et al., 2007; Gaver, 2012). Our research had three phases: ethnographic observations (Phase 1), a poster-based discussion with diverse stakeholders (Phase 2), and a prototype deployment (Phase 3). This methodological approach was chosen because our research focuses on a problem that does not have a single optimal solution and requires the generation of new design knowledge through iterative cycles of making and reflecting.
@@
3.1. Participants
For Phase 1, we collaborated with a local nursing home to recruit 15 participants for our ethnographic observations. We then conducted semi-structured interviews with 12 of these participants after the ethnographic observations. For Phase 2, we recruited 9 stakeholders. Three were researchers specializing in geriatric medicine, public health, and biomedical sciences, another three were interaction designers with experience in healthcare and systems for older adults, and the remaining three were older adults who represented our target end users. These stakeholders had diverse perspectives and levels of prior engagement with personal data, clinical, design, and lived-experiences, similar to prior work (Karyda et al., 2021). For Phase 3, we recruited 5 older adults with sedentary habits. Each participant hosted the CushLumi systemin their home for a full day and then participated in semi-structured interviews to capture their lived experience using the system. Herein, we refer to older adult participants as P1–P15, expert researchers as R1–R3, and interaction designers as D1–D3 (Table 1).
@@ 3.2. Data Physicalization Design Concepts and the CushLumi Prototype
Insights from Phase 1 led us away from a smartphone‑based intervention toward feedback that could be quietly embedded in everyday surroundings. Participants reported disliking complex apps and “nagging” notifications, and cushions repeatedly appeared as mundane but central artefacts in long sitting episodes. At the same time, prior work highlights cushions and lighting as familiar, pervasive elements of older adults’ domestic interiors and as promising channels for low‑priority ambient feedback (Wisneski et al., 1998; Davis et al., 2017a; McLaughlin and Pak, 2020; Feine, 2025). Building on this, we developed CushLumi as a two‑part, hybrid data physicalization system: a sensing cushion that captures sedentary behaviour, and a separate ambient lamp that materializes this data in the home.
Figure 2 presents an overall schematic of the CushLumi system, while Figure 3, footnote 2 and Figure 5 illustrate its layered architecture, hardware components, and light behaviour. In everyday use, the cushion detects when an older adult sits down, stands up, and remains seated for prolonged periods. These events are processed in a simple cloud‑based pipeline that tracks sit durations and drives a small wooden lamp placed nearby. As uninterrupted sitting accumulates, the lamp’s colour and brightness gradually shift from a soft warm white toward a more intense, reddish glow; when the person stands, the lamp briefly plays a colourful celebratory animation before returning to its baseline state. Laser‑cut patterns on the lamp depict positive everyday activities (e.g., stretching, gardening), and a touch‑sensitive top surface serves as a mute button, allowing users to temporarily silence the feedback and thereby maintain a sense of agency. Further design explorations and rationale via RtD, technical details regarding the sensing hardware, light‑mapping logic, and cloud infrastructure are provided in Appendix A and Appendix B.
@@ 3.3. Procedures
Herein, we describe the procedures for each phase of our RtD methodology.
@@ 3.3.1. Phase 1
To understand the context of sedentary behavior and older adults’ relationship with technology, we first conducted ethnographic observations of 15 older adults in a local nursing home. We focused on their daily routines, sitting habits, and interactions with everyday objects and technologies. Field notes documented locations, durations, and social contexts of sitting, as well as observed uses of digital devices. Following the observations, we conducted semi-structured interviews with 12 of these older adults. 3 participants were excluded from the interview phase due to scheduling conflicts or health-related fatigue. The interviews covered topics such as daily activity patterns, attitudes toward sedentary behavior, experiences with digital technologies (including smartphones and other screen-based devices), and preferences for how health-related information should be communicated. Interviews also invited participants to reflect on specific situations observed during the ethnographic phase.
@@
Figure 4. The components of the CushLumi system. The CushLumi system major used Particle Photon microcontrollers [^1] Particle Photon microcontrollers: https://www.particle.io/, which was programmed in C++, to process real-time data from the flex wiring sensors and control the LED circle. The wireless communication between the hardware components was controlled using Particle’s cloud platform. IFTTT [^2] IFTTT: https://ifttt.com/ was also used to prototype simple time-based and remote control of the system (e.g., muting or unmuting the light via a phone widget or a scheduled applet).
@@ 3.3.3. Phase 3
We obtained informed consent from all participants before each phase. For older adults in the nursing home and in-home deployments, the study procedures were also coordinated with facility staff and family members when appropriate. Participants could withdraw at any time without penalty, and all collected data was pseudonymized during analysis and for reporting herein.
@@ 3.4. Data Collection and Analysis
We analyzed the semi-structured interview data using reflexive thematic analysis following Braun and Clarke’s approach, as adapted in prior HCI research (Lin, 2019; Bowman et al., 2023). Two researchers independently open-coded the transcripts, identifying segments related to perceptions of sedentary behavior, reactions to different forms of data representation, emotional responses, and experiences with the prototypes. They then met to discuss and merge their findings into a unified codebook. This codebook was systematically applied to the entire dataset, with the researchers working to refine the codes and group them into higher-level themes. The resulting themes were then related back to the three phases of the RtD process, informing both the iterative design decisions described above and the final set of findings reported herein. Note that we use annotations such as P3_\_Phase1 to indicate that, in this example, a quote was from P3 during Phase 1.